# MedMamba: Vision Mamba for Medical Image Classification

### v4

@@ Abstract
Since the era of deep learning, convolutional neural networks (CNNs) and vision transformers (ViTs) have been extensively studied and widely used in medical image classification tasks. Unfortunately, CNN’s limitations in modeling long-range dependencies result in poor classification performances. In contrast, ViTs are hampered by the quadratic computational complexity of their self-attention mechanism, making them difficult to deploy in real-world settings with limited computational resources. Recent studies have shown that state space models (SSMs) represented by Mamba can effectively model long-range dependencies while maintaining linear computational complexity. Inspired by it, we proposed MedMamba, the first vision Mamba for generalized medical image classification. Concretely, we introduced a novel hybrid basic block named SS-Conv-SSM, which integrates the convolutional layers for extracting local features with the abilities of SSM to capture long-range dependencies, aiming to model medical images from different image modalities efficiently. By employing the grouped convolution strategy and channel-shuffle operation, MedMamba successfully provides fewer model parameters and a lower computational burden for efficient applications. To demonstrate the potential of MedMamba, we conducted extensive experiments using 16 datasets containing ten imaging modalities and 411,007 images. Experimental results show that the proposed MedMamba demonstrates competitive performance in classifying various medical images compared with the state-of-the-art methods. Our work is aims to establish a new baseline for medical image classification and provide valuable insights for developing more powerful SSM-based artificial intelligence algorithms and application systems in the medical field. The source codes and all pre-trained weights of MedMamba are available at https://github.com/YubiaoYue/MedMamba.
@@ 3.2.1 Overall architecture
MedMamba drew inspiration from VMamba’s architecture design and ideas. Figure 1 shows the overall architecture of MedMamba. Specifically, MedMamba includes a patch embedding layer, stacked SS-Conv-SSM blocks, patch merging layers for down-sampling, and a feature classifier.
@@
The 2D-selective-scan (SS2D) (Figure 2) proposed by VMamba is one of the core elements of MedMamba. SS2D inherits the selective scan space state sequence model (S6) designed for natural language processing tasks and successfully solves the "direction-sensitive" problem in S6. Specifically, to narrow the gap between 1-D array scanning and 2-D plain traversing, SS2D introduced a Cross-Scan Module (CSM), facilitating the extension of S6 to visual data without affecting the reception field. CSM adopts a four-way scanning strategy, i.e., from four corners all across the feature map to the opposite location to traverse the spatial domain of image feature maps, which ensures that each pixel in a feature map integrates information from all other locations in different directions, resulting in a global receptive field without increasing the linear computational complexity.
By integrating CSM, SS2D inherits the linear complexity of S6 while capturing long-range dependencies, which is essential for achieving accurate medical image classification. The SS2D contains three components consists: a scan expanding operation (CSM), an S6 block, and a scan merging operation. Figure2 visually presents the internal mechanism of SS2D. Concretely, the scan expanding operation first unfolds the input image along four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right) into sequences. An S6 block then processes all sequences to extract the features, ensuring that information from various directions is thoroughly scanned. Lastly, the four output features from the four directions are merged through scan merging to construct the final 2D feature map, resulting in a final output of the same size as the input. The S6 block originates from Mamba and introduces a selective mechanism based on S4 by adjusting the parameters of SSM according to input. This enables the model to distinguish and retain relevant information while filtering out irrelevant information. The pseudo-code for the S6 block is presented in Algorithm 1.

### v5

@@ Abstract
Since the era of deep learning, convolutional neural networks (CNNs) and vision transformers (ViTs) have been extensively studied and widely used in medical image classification tasks. Unfortunately, CNN’s limitations in modeling long-range dependencies result in poor classification performances. In contrast, ViTs are hampered by the quadratic computational complexity of their self-attention mechanism, making them difficult to deploy in real-world settings with limited computational resources. Recent studies have shown that state space models (SSMs) represented by Mamba can effectively model long-range dependencies while maintaining linear computational complexity. Inspired by it, we proposed MedMamba, the first Vision Mamba for generalized medical image classification. Concretely, we introduced a novel hybrid basic block named SS-Conv-SSM, which purely integrates the convolutional layers for extracting local features with the abilities of SSM to capture long-range dependencies, aiming to model medical images from different image modalities efficiently. By employing the grouped convolution strategy and channel-shuffle operation, MedMamba successfully provides fewer model parameters and a lower computational burden for efficient applications without sacrificing accuracy. We thoroughly evaluated MedMamba using 16 datasets containing ten imaging modalities and 411,007 images. Experimental results show that MedMamba demonstrates competitive performance on most tasks compared with the state-of-the-art methods. This work aims to explore the potential of Vision Mamba and establish a new baseline for medical image classification, thereby providing valuable insights for developing more powerful Mamba-based artificial intelligence algorithms and applications in medicine. The source codes and all pre-trained weights of MedMamba are available at https://github.com/YubiaoYue/MedMamba.
@@ 3.2.1 Overall architecture
Figure 1 shows the overall architecture of MedMamba. Specifically, MedMamba includes a patch embedding layer, stacked SS-Conv-SSM blocks, patch merging layers for down-sampling, and a feature classifier.
@@
The 2D-selective-scan (SS2D) (Figure 2) proposed by VMamba[42] is one of the core elements of MedMamba. SS2D inherits the selective scan space state sequence model (S6) designed for natural language processing tasks and successfully solves the "direction-sensitive" problem in S6. Specifically, to narrow the gap between 1-D array scanning and 2-D plain traversing, SS2D introduced a Cross-Scan Module (CSM), facilitating the extension of S6 to visual data without affecting the reception field. CSM adopts a four-way scanning strategy, i.e., from four corners all across the feature map to the opposite location to traverse the spatial domain of image feature maps, which ensures that each pixel in a feature map integrates information from all other locations in different directions, resulting in a global receptive field without increasing the linear computational complexity.
By integrating CSM, SS2D inherits the linear complexity of S6 while capturing long-range dependencies, which is essential for achieving accurate medical image classification. The SS2D contains three components consists: a scan expanding operation (CSM), an S6 block, and a scan merging operation. Figure2 visually presents the internal mechanism of SS2D. Concretely, the scan expanding operation first unfolds the input image along four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right) into sequences. An S6 block then processes all sequences to extract the features, ensuring that information from various directions is thoroughly scanned. Lastly, the four output features from the four directions are merged through scan merging to construct the final 2D feature map, resulting in a final output of the same size as the input.
The S6 block originates from Mamba and introduces a selective mechanism based on S4 by adjusting the parameters of SSM according to input. This enables the model to distinguish and retain relevant information while filtering out irrelevant information. The pseudo-code for the S6 block is presented in Algorithm 1.
For ease of understanding, we denoted the input of SS2D, i.e., the feature map, as $I\in{R^{H\times W\times C}}$, where $I$[$h$][$w$] represents one token in the h-th row and w-th column of the feature map. Therefore, the scan expanding operation can be formalized as follows[87]:
${s_{1}}[i]=I\left[{i\bmod W}\right]\left[{\left\lfloor{i/W}\right\rfloor}\right]$ (5)
${s_{2}}[i]=I\left[{(N-i-1)\bmod W}\right]\left[{\left\lfloor{(N-i-1)/W}\right\rfloor}\right]$ (6)
${s_{3}}[i]=I\left[{\left\lfloor{i/H}\right\rfloor}\right]\left[{\left\lfloor{i\bmod H}\right\rfloor}\right]$ (7)
${s_{4}}[i]=I\left[{\left\lfloor{(N-i-1)/H}\right\rfloor}\right]\left[{(N-i-1)\bmod H}\right]$ (8)
Where $N=H\times W$, $0\leq i<N$, ${s_{1}},{s_{2}},{s_{3}},{s_{4}}\in{R^{N\times C}}$ are the expanded and flattened token sequences. Next, we employed the S6 to selectively scan each token sequence, resulting in contextual token sequences ${s^{\prime}_{1}},{s^{\prime}_{2}},{s^{\prime}_{3}},{s^{\prime}_{4}}\in{R^{N\times C}}$. Immediately after, SS2D apply reversed operations to the contextual token sequences:
${I^{\prime}_{1}}[i][j]={s^{\prime}_{1}}[j\times W+i]$ (9)
${I^{\prime}_{2}}[i][j]={s^{\prime}_{2}}[N-1-j\times W-i]$ (10)
${I^{\prime}_{3}}[i][j]={s^{\prime}_{3}}[i\times H+j]$ (11)
${I^{\prime}_{4}}[i][j]={s^{\prime}_{4}}[N-1-i\times H-j]$ (12)
Where ${I^{\prime}_{1}},{I^{\prime}_{2}},{I^{\prime}_{3}},{I^{\prime}_{4}}\in{R^{H\times W\times C}}$ denote the expanded and transformed feature map of I. In the end, SS2D applies the scan merging to obtain the output:
$I^{\prime}={I^{\prime}_{1}}+{I^{\prime}_{2}}+{I^{\prime}_{3}}+{I^{\prime}_{4}}$ (13)